Because of the pandemic, no observers can come to the telescope. This pool is entirely in remote mode.
A couple of preparatory notes by Stefano:
- please use the PIIC monitors. If you notice that one (or more) of the PIIC monitors stop processing the data, please notify the AoD: it is possible that the scripts that produce the IMBFITS files have stopped.
- please take numerous calibrators scans, also in bad sky conditions (as long as NIKA2 is used): for example take one calib scan whenever the pointing/focus source is a calibrator. This will help the analysis of the calibration. Thank you in advance.
Little extra by Stefano: I would have liked to be at the telescope to celebrate the 50th pool of NIKA2 with a special mid-night pasta, but unfortunately it is still not possible. Therefore I have prepared a little surprise. I hope it can make you smile in these eerie times. Stay Safe!
Feb 9th, Tuesday
16h - 21h; Angel, Bilal: Stopped. Bad weather. Not a single scan.
21h - 03h; Carsten: Stopped.
Feb 10th, Wednesday
03h - 09h; Stopped.
09h - 15h; Still stopped.
15h - 21h; Stopped.
21h - 03h; Stopped.
Feb 11th, Thursday
03h - 09h; Sergio (Op.), Angel: Checking the telescope at 05h (LT). Opacity ~0.8. NIKA2 Pool targets according to these conditions are too low. Switching to EMIR backup project 171-20. Pointing/focus OK. Wobbler required, but it doesn't work after several tries. Switching to NIKA2 at ~7 am. Resonances reloaded. Opacity ~0.65. Pointing/focus on K3-20A. Despite the opacity level, conditions look quite stable. Trying to restart the data-filler of TAPAS; it doesn't work (waiting for Computer Group instructions). NGC7027: calib_1scan (s32). Then, a skydip. Monitoring stability; large spikes of opacity.
09h - 15h; Antonio, Emmanuel, Bilal: stopped ~1h for technical tests of wobbler system. Observed with NIKA2 for project 126-20.
15h - 21h; Angel: Observed 12 fields of 126-20, and started with EMIR backup program 109-20 at 20h (LT); tau~0.7, unstable. Data records still being recovered in TAPAS.
21h - 03h; Annie, Carsten take over to continue with EMIR projects, first 109-20. Tau~0.7. Some clouds. At UT00:00 change to EMIR 097-20 154.05+5.07. Tau drops to 0.5. TAPAS working.
Feb 12th, Friday
03h - 09h; Miguel, Angel: Starting with opacity ~0.8; large instabilities. Continuing with EMIR backup programs: 171-20. Four bright targets observed. The opacity at the end of the shift dropped quickly to ~0.3. Switching to NIKA2 at 08:45. Resonances reloaded after that. TAPAS takes more time than usual to update the records: reported to IRAM's Granada User Support.
09h - 15h; Laurence, Miren: Starting with opacity at ~0.3, increasing. Focus on NGC7027. Difficult to get a good pointing, clear indication of atmospheric instabilities. Whenever we obtained "reasonable" pointing, we did calib_1scan scans on the three secondary calibrators (MWC349, NGC7027 and CRL2688). The quality of these observation is uncertain. To be repeated. At ~ 11h (LT) start with 126-20 project, one repetition toward 4 different GP fields. Opacity drops and after focus and pointing, we change to 199-16 and performed 2 repetitions on PSZ2G104 (dubious atmospheric stability). We probed atmospheric condition toward the opposite azimuth: better focus and pointing. We moved back to 199-16 but choose another target close to this azimuth: 3 repetitions on ACTJ2302. The atmosphere was stable at the end of our shift.
15h - 21h; Isabelle, Karine: At the begining of the shift, tau was very good (~0.15) and the atmosphere stable. We started with a focus, pointing, beammap, and a calib_1scan on Uranus. We then observed project 094-20 (5 Nrep on L1506). We then did a skydip and continued on project 094-20 observing G150 (5 NRep). Tau was ~ 0.2 by the end of the shift and the atmosphere stable all over the shift.
21h - 03h; Golshan, Carsten: 160-16: NGC3198 (2h) then NGC5194 (2h), tau~0.2-0.28, stable with <+/-6Jy, rms~2Jy. Pointing, Focus stable. 1 skydip. Tau prediction page failed (saying 502 Bad Gateway)
Feb 13th, Saturday
03h - 09h; Ankur, Kaustuv: We started with tau ~ 0.18 and it remained stable throughout the shift. Atmospheric conditions were very good. We did 2 repetitions on 180-20 project and then did 7 repetitions on PSZ2G104 (199-16). It was a smooth shift.
09h - 15h; Antonio, Marco: The opacity has been always around 0.10-0.15. We did 5 repetitions on PSZ2-G087 (12/12 completed! - proj (199-16)), 2 repetitions on PSZ2-G045 (2/6 - proj (199-16)), 1 skydip and 2 repetitions on G82 (proj (094-20)).
15h - 21h; Angel: Starting with tau~0.2. Project 094-20: 2 rep. on G82, 2 rep. on G155, and 1 rep. on G150 (finished). Note of CK: 094-20 scan 155 shows stability +/-10Jy, rms=4.4Jy @2mm. Calibrations: beam-map (h) and a calib_1scan on Uranus (mean tau~0.27). Project 138-20: 1 rep. on NGC2701 (tau~0.2 and very stable at the beginning).
21h - 03h; Carsten: 160-16 NGC3198 tau~0.2, Atmosphere not stable, slowly improving from rms before/after sky subtraction: 8-19Jy/b(!!) / 40mJy/b to 3Jy/b / 33mJy/b. Tau also slowly improving to 0.14. At midnight local time changing to 180-20 with stable atmosphere (rms=1Jy/b).
Feb 14th, Sunday
03h - 09h; Ankur, Kaustuv: We started with tau ~0.17 and it remained stable throughout the shift. Atmospheric conditions were also stable. We did 2 repetitions on 180-20 project and then did 9 repetitions on PSZ2G104 (project 199-16; Total target Nreps required for PSZ2G104 - 16/16 completed).
11h UT NOTE BY STEFANO: the make-imbfits for Arrays 1 and 3 seems to have stopped around 00:30 of Feb 13th. The situation of pool 49 is repeating. Please, start it again! I also urge the users to check this regularly and use the PIIC monitor for pointing (Ar2) and focus (Ar1).
09h - 15h; Aishwarya & Laurence: tau225 around 0.15 and stable atmosphere (no hint of anomalous refraction from the pointing). We observed 199-16 (PSZ2G108, PSZ2G045 and ACTJ2302) and did one high-quality calib_1scan on MWC349. The beam started to enlarge a bit at the end of our shift.
15h - 21h; Matthieu: The weather conditions were excellent with tau in the 0.15-0.18 range. We started with a beammap and a skydip. We then did 3 repeats of G157 and then 2 repeats of G93 (094-20). We then switched to 118-20 (SH2), when it was bright enough and completed this program. Everything went smoothly. Note of CK: Scan 203 of 094-20 shows excellent stability: 0.9Jy/b rms at 2mm.
21h - 03h; Carsten: finishing last scan on 118-20. Tau~0.17-0.21: Concerning 1mm arrays: IMBfits is working on scan 154 of yesterday, still far from scan 231 of what is needed now. Angel restarts the acquisition. Fraction of healthy KIDs increased. calib_1scan & beam map on 3C84. 160-16: NGC2976 (2h), M51 (2h, tau=0.22-0.24). Exceptionally stable conditions: 1Jy/b rms @ 2mm.
Feb 15th, Monday
03h - 09h; Guilaine: Tau between 0.23 and 0.28, but most of the time around 0.23-0.24. Atmosphere quite stable! Observations 4.21 hours on GOODSN (N2CLS). Then switch to PSZ2G108 (LPSZ) after a pointing/focus session from 8:15 to 8:30 (fwhms are very good, atm very stable).
09h - 15h; Emmanuel & Frédéric : tau225 around 0.2 and stable atmosphere. We observed 199-16 (PSZ2G108 and ACTJ2302) and did one calib_1scan on NGC7027. The beam started to enlarge a lot at the end of our shift and hence we have not done any beammap.
15h - 21h; Angel: tau255 ~0.25; dish relaxing; pointing & focus on 3C84. Skydip. Project 094-20: 0.22<tau255< ; target G155 x 3.5 reps. Very stable at the beginning: s167, +/-2 Jy/beam p2p, <rms>~0.8 Jy/beam (2 mm). Project 093-20: , target NGC2071 x 2 reps. Stability, 20:50 LT: on a 20' timeline, +/-5 Jy/beam, rms~1 Jy/beam (2 mm). Note of Bilal: weather.iram.es is back on.
21h - 03h; Golshan, Annie, Carsten: tau~0.25-0.27: 138-20: 1 scan on ngc2071. Beam map & calib_1scan on 3C84. M51: 4 hours (stable conditions: < +/-4Jy/b @ 2mm).
Feb 16th, Tuesday
03h - 10:30; Xavier: Smooth observations. Opacity went down from 0.2 to 0.15. Did some repeats on 192-16/Goodsnorth (3 to 6 LT) then moved on 199-16 to finish PSZG2G108 with 7 repeats. Did one Goodsnorth again to finish the shift with an excellent beam. Conditions degraded (beamwise) at 10 LT.